OpenStudy (anonymous):

Express 108 as a product of primes.

OpenStudy (solomonzelman):

\[108=2 \times 2 \times 3 \times 3 \times 3\]

OpenStudy (solomonzelman):

\[108= 4 \times 27=2^2 \times 3^3=2 \times 2 \times 3 \times 3 \times 3\]